OpenStudy (anonymous):

**MEDAL + FAN** Devin throws a football 7 1/3 yards. Carl throws it 2 1/2 times farther. How much farther did Carl's throw travel than Devin's? A. 2 1/2 yards B. 3 1/3 yards C. 11 yards <--- D. 18 1/3 yards Formula? @gorv

OpenStudy (gorv):

7 1/3=22/3

OpenStudy (gorv):

2 1/2=5/2 so5/2 times

OpenStudy (gorv):

\[=\frac{ 22 }{ 3 }*\frac{ 5 }{ 2 }=\frac{ 55 }{ 3 }\]

OpenStudy (asevilla5):

Devin throws 7 1/3 = 22/3 yards Carl throws 2 1/2 times = 5/2 times farther So Carl throws 22/3 * 5/2 = 55/3 yards 55/3 yards - 22/3 yards = 33/3 yards = 11 yards

OpenStudy (anonymous):

Thanks! :) I'll fan you soon once my client stops bugging out.